OpenStudy (sleepyjess):

Question below @ganeshie8

OpenStudy (sleepyjess):

Solve the triangle. A = 32, a = 19, b = 14

OpenStudy (unklerhaukus):

The law of sines says: \[\frac{\sin A}a=\frac{\sin B}b=\frac{\sin C}c\]

ganeshie8 (ganeshie8):

you may use the first proportion to solve angle B \[\frac{\sin A}a=\frac{\sin B}b\]

OpenStudy (sleepyjess):

sin(32)/19=sin(x)/14 sin(32)19 * 14 = sin(x)

ganeshie8 (ganeshie8):

do you mean (sin(32)\(\color{red}{/}\)19) * 14 = sin(x)

OpenStudy (sleepyjess):

Oh, yes :)

ganeshie8 (ganeshie8):

solve \(x\)

ganeshie8 (ganeshie8):

i see you mean \(x = B\)

OpenStudy (freckles):

you might have to check for two triangles

ganeshie8 (ganeshie8):

@ziqbal103 your questions are a bit tough.. :/

OpenStudy (anonymous):

when i tag you you don't rply me @ganeshie8

OpenStudy (anonymous):

in my posting questions

OpenStudy (sleepyjess):

Thank you for the help @ganeshie8 :)